Benjamin and associates, A real estate developer recently built 210 condos. The condos are either two bedroom units or three bedroom units . If the total number of rooms in the entire complex is 535 how many two bedroom units are there? How many three-bedroom?
Answer by macston(5194)

(If you mean there are 535 bedrooms total)
.
x=number of two bedroom; y=number of three bedroom
.
x+y=210
x=210-y Use this to substitute for x
.
2x+3y=535 Substitute for x from above
2(210-y)+3y=535
420-2y+3y=535 Subtract 420 from each side
y=115
ANSWER 1: There were 115 three bedroom condos
.
x=210-y
x=210-115
x=95
ANSWER 2: There were 95 two bedroom condos
.
CHECK:
2x+3y=535
2(95)+3(115)=535
190+345=535
535=535
.